OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3?

OpenStudy (lgbasallote):

use \[y - y_1 = m(x-x_1)\] substitute \[y - 6 = 3(x - 1)\]

OpenStudy (lgbasallote):

can you simplify that now?

OpenStudy (anonymous):

No.

OpenStudy (lgbasallote):

distribute 3 into x - 1 what do you get

OpenStudy (anonymous):

y = 3x + 3?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Awesome Thanks =)